OpenStudy (anonymous):

Simplfy (n) - (n-1) (r) (r)

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ r\end{matrix}\right) - \left(\begin{matrix}n-1 \\ r\end{matrix}\right)\]

OpenStudy (anonymous):

simplify

OpenStudy (welshfella):